graylove:

a+d+c=4 a2+b2+c2=10 a3+b3+c3=22 a4+b4+c4=?

Alexis1415:

Are these answer choices or a question?

graylove:

its a question

Alexis1415:

So do you know what you are supposed to do with the numbers? Are you supposed to solve them?

graylove:

no and you have to write it out

YTStresssPG3D:

@alexis1415 wrote:
So do you know what you are supposed to do with the numbers? Are you supposed to solve them?
im confused to

graylove:

yes same like what do you mean by that

mhanifa:

Can you express the 2nd, 3rd and 4th powers of the sum (a + b + c)?

YTStresssPG3D:

are you trying to solve the last one?

Alexis1415:

Oh i think you have to try to solve the last one like YTStresss said

graylove:

yes im trying to solve that last one

Alexis1415:

okay i will help one second i have to write it out

YTStresssPG3D:

ok

graylove:

no i need the the last one solved ok

YTStresssPG3D:

so what do you think it is

Alexis1415:

First: a+b+c=4 Squaring both sides, a^2+b^2+c^2 + 2(ab + bc + ca) = 16 10+2(ab+bc+ca)=16 ab + bc + ca=3

graylove:

oh ok thanks?

Alexis1415:

Second: a^3+b^3+c^3 - 3abc = (a + b + c) (a^2+b^2+c^2 - (ab + bc + ca)) 22-3abc=(4)(10-3) 22-3abc=28 3abc=-6 abc= -2

graylove:

ok

Alexis1415:

Next: ab + bc + ca=3 (ab)^2 + (bc)^2 + (ac)^2 + 2(ab.bc + ab.ca + bc.ca) = 9) (ab)^2 + (bc)^2 + (ac)^2 + 2abc(a + b + c) = 9 (ab)^2 + (bc)^2 + (ac)^2 + 2(-2)(4) = 9 (ab)^2 + (bc)^2 + (ac)^2 = 25 a^2b^2 + b^2c^2 + a^2c^2 = 25

graylove:

ok

graylove:

well i have to go rn so i will see the chat later

Alexis1415:

Last: a^2+b^2+c^2=10 a^4 + b^4 + c^4 + 2( a^2b^2 + b^2c^2 + a^2c^2) = 100 a^4 + b^4 + c^4 + 2( 25) = 100 a^4 + b^4 + c^4 + 50 = 100 a^4 + b^4 + c^4 = 50

graylove:

oh ok